F or the follow ing differential equation locate and classify its singular points on the x-axis


𝑥²𝑦′′+ (2−𝑥)𝑦′= 0.

1
Expert's answer
2021-12-21T17:46:06-0500

"y''+\frac{2-x}{x^2}y'=0"


"P(x)=\frac{2-x}{x^2}"


x = 0 is singular point

"(x-0)P(x)=\frac{2-x}{x}" is not analytic at x = 0

so, x = 0 is irregular singular point


x = 2 is singular point

"(x-2)P(x)=-\frac{(x-2)^2}{x^2}" ​is analytic at x = 2

so, x = 2 is regular singular point